OpenStudy (anonymous):

check answer again solve 2x^2=6

OpenStudy (anonymous):

find the x intercept of f(x)=2x^2-6

OpenStudy (anonymous):

factor: 2(x^2-3), set = to 0 x^2-3=0 x^2=3, take the square root of both sides x=+ or - \[\sqrt{3}\]

OpenStudy (anonymous):

i got √3, -√3 for the first part (0,-6)

OpenStudy (anonymous):

The y intercept is (0,-6

OpenStudy (anonymous):

x intercept set y = 0. y -intercept set x = 0

OpenStudy (anonymous):

(0,-6)

OpenStudy (anonymous):

what lagrangeson said

OpenStudy (anonymous):

i have what I came up with is that right

OpenStudy (anonymous):

yes, you got

OpenStudy (anonymous):

thx